# Perception, Reason, Think, and Plan: A Survey on Large Multimodal Reasoning Models

### v1

@@ 5.2.2 Joint Multimodal Generation
The development of multimodal large language models (MLLMs) has significantly advanced interleaved image-text generation, with datasets like MM-Interleaved (Tian et al., 2024) and ANOLE (Chern et al., 2024) supporting model training with high-quality annotated image-text pairs. These datasets emphasize the need for models to generate contextually relevant and visually coherent content. Benchmarks like InterleavedEval (Liu et al., 2024c) and OpenLEAF (An et al., 2024) focus on evaluating models’ ability to generate coherent and aligned image-text pairs, while OpenING (Zhou et al., 2024a) provides a more diverse set of tasks to assess interleaved image-text generation.
@@ 5.3.1 General Visual Reasoning
General visual reasoning is one of the most critical capabilities in Multimodal Reasoning Models. It requires models not only to perceive visual information but also to comprehend, analyze, and reason about it using extensive knowledge, logical deduction, and common sense across a variety of scenarios.
To rigorously assess this ability, a wide range of benchmarks has been developed, each targeting distinct aspects of visual reasoning. Moving beyond simple question answering tasks (e.g., VQA), Visual Commonsense Reasoning benchmarks like VCR (Zellers et al., 2019), and specialized datasets like PhysBench (Chow et al., 2025) for physical reasoning, and VideoPhy (Bansal et al., 2024) for understanding physical common sense in videos, challenge models to apply everyday knowledge to interpret visual situations.
@@ 5.3.2 Domain-specific Reasoning
Domain-specific reasoning benchmarks play a crucial role in evaluating the specialized reasoning capabilities of multimodal models in specific fields. For mathematical reasoning, datasets like MathVista (Lu et al., 2024) and MATH-Vision (Wang et al., 2024c) assess a model’s ability to solve mathematical problems in visual contexts, requiring both visual understanding and mathematical inference. Similarly, benchmarks like ChartQA (Masry et al., 2022) and ScienceQA (Lu et al., 2022) focus on reasoning in specific domains.
@@
For physical reasoning, datasets like PhysBench (Chow et al., 2025), VideoPhy (Bansal et al., 2024), and CRAVE (Sun et al., 2025) assess models’ understanding of physical laws and common sense across visual and video contexts. Finally, benchmarks like GAIA-1 (Hu et al., 2023) and RoboGen (Wang et al., 2024l) support the development of world models by evaluating how well models can simulate and reason about real-world dynamics and interactions.
@@ 5.4.1 GUI Navigation
Benchmarks in GUI navigation assess agents’ abilities to plan and execute tasks across digital interfaces, requiring robust visual-language grounding and multi-step reasoning. WebArena (Zhou et al., 2024c) and Mind2Web (Deng et al., 2023) offer realistic web environments for navigation and information extraction, with Mind2Web further introducing cross-website tasks to test generalizability. VisualWebBench (Liu et al., 2024b) advances visual-intensive planning with 1.5K tasks focused on cross-page integration and element localization. Windows Agent Arena (Bonatti et al., 2024) evaluates cross-application planning in desktop environments, while Ferret-UI (You et al., 2024) focuses on grounded UI understanding for executing multi-step instructions. Benchmarks like WebShop (Yao et al., 2022) test visual-language grounding in simulated e-commerce environments. Similarly, OSWorld (Xie et al., 2024a) and OmniACT (Kapoor et al., 2024) provide real desktop OS environments, supporting cross-application workflows such as file manipulation and data processing. VisualAgentBench (Liu et al., 2024d) extends this paradigm by systematically evaluating large multimodal models across GUI, embodied, and visual design tasks, establishing a unified benchmark for planning and acting in visually rich digital environments. This is complemented by benchmarks like LlamaTouch (Zhang et al., 2024d), which scales mobile UI automation with 495 tasks, testing multi-step operations such as app navigation.
@@ 5.5 Evaluation Method
Due to the diversity of answers, exact and fuzzy matching methods are often unable to encompass all candidate options. To ensure fairness and accuracy in evaluation, the Options Matching approach was introduced. In this method, the system prompt includes several candidate options, and the model is required to select the most appropriate one. Moreover, to reduce the possibility of the model exhibiting a preference for a specific option during the selection process, works such as MMBench (Liu et al., 2024f) have adopted the CircularEval methodology to minimize stochastic variations in the evaluation.
@@
During the evaluation process, the capabilities of a single model are inherently limited, which may lead to shortcomings when processing diverse multimodal information. To this end, agent-based approaches can leverage tools to mitigate the inherent limitations of the model itself. For instance, CIGEval (Wang et al., 2025f) expands the visual understanding capabilities of MLLMs by integrating a multi-functional toolbox, thereby enabling more fine-grained evaluation. Moreover, multi-agent discussions have shown effectiveness in downstream tasks by fostering consensus and producing more robust solutions (Xu et al., 2023b; Chen et al., 2024d; Xu et al., 2025f), a benefit that also extends to evaluation settings. Methods that leverage cooperative or adversarial interactions among multiple agents to assess outputs have demonstrated more reliable and interpretable evaluations (Chan et al., 2024; Li et al., 2024h; Zhao et al., 2024b; Liang et al., 2024).
@@ 6 Conclusion
In this paper, we survey the evolution of multimodal reasoning models, highlighting pivotal advancements and paradigm-shifting milestones in the field. While current models predominantly adopt a␣language-centric reasoning paradigm—delivering impressive results in tasks like visual question answering, visual math, and video understanding—critical challenges persist. Notably,␣visual-centric long reasoning (e.g., understanding 3D contexts, addressing complex visual information-seeking questions) and␣interactive multimodal reasoning (e.g., dynamic cross-modal dialogue or iterative feedback loops) remain underdeveloped frontiers requiring deeper exploration.
Building on empirical evaluations and experimental insights, we propose a forward-looking concept for␣inherently multimodal large models that transcend language-dominated architectures. Such models should prioritize three core capabilities: Multimodal Agentic Reasoning: Enabling proactive environmental interaction (e.g., embodied AI agents that learn through real-world trial and error). Omini-Modal Understanding and Generative Reasoning: Integrating any-modal semantics (e.g., aligning abstract concepts across vision, audio, and text) while resolving ambiguities in complex, open-world contexts; Producing coherent, context-aware outputs across modalities (e.g., generating diagrams from spoken instructions or synthesizing video narratives from text). By addressing these dimensions, future models could achieve human-like contextual adaptability, bridging the gap between isolated task performance and generalized, real-world problem-solving.

### v2

@@ 5.2.2 Joint Multimodal Generation
The development of multimodal large language models (MLLMs) has significantly advanced interleaved image-text generation, with datasets like MM-Interleaved (Tian et al., 2024) and ANOLE (Chern et al., 2024) supporting model training with high-quality annotated image-text pairs. These datasets emphasise the need for models to generate contextually relevant and visually coherent content. Benchmarks like InterleavedEval (Liu et al., 2024c) and OpenLEAF (An et al., 2024) focus on evaluating models’ ability to generate coherent and aligned image-text pairs, while OpenING (Zhou et al., 2024a) provides a more diverse set of tasks to assess interleaved image-text generation.
@@ 5.3.1 General Visual Reasoning
General visual reasoning is one of the most critical capabilities in Multimodal Reasoning Models. It requires models not only to perceive visual information but also to comprehend, analyse, and reason about it using extensive knowledge, logical deduction, and common sense across a variety of scenarios.
To rigorously assess this ability, a wide range of benchmarks has been developed, each targeting distinct aspects of visual reasoning. Moving beyond simple question answering tasks (e.g., VQA), Visual Commonsense Reasoning benchmarks like VCR (Zellers et al., 2019), and specialised datasets like PhysBench (Chow et al., 2025b) for physical reasoning, and VideoPhy (Bansal et al., 2024) for understanding physical common sense in videos, challenge models to apply everyday knowledge to interpret visual situations.
@@ 5.3.2 Domain-specific Reasoning
Domain-specific reasoning benchmarks play a crucial role in evaluating the specialized reasoning capabilities of multimodal models in specific fields. For mathematical reasoning, datasets like MathVista (Lu et al., 2024) and MATH-Vision (Wang et al., 2024c) assess a model’s ability to solve mathematical problems in visual contexts, requiring both visual understanding and mathematical inference. Similarly, benchmarks like ChartQA (Masry et al., 2022), ScienceQA (Lu et al., 2022) and CharXiv (Wang et al., 2024n) focus on reasoning in specific domains.
@@
For physical reasoning, datasets like PhysBench (Chow et al., 2025b), VideoPhy (Bansal et al., 2024), and CRAVE (Sun et al., 2025) assess models’ understanding of physical laws and common sense across visual and video contexts. Finally, benchmarks like GAIA-1 (Hu et al., 2023) and RoboGen (Wang et al., 2024l) support the development of world models by evaluating how well models can simulate and reason about real-world dynamics and interactions.
Several datasets target reasoning in specific high-value domains. WeThink-Dataset contributes a richly annotated dataset with explicit reasoning paths, enabling instruction tuning and reinforcement learning for vision-centric reasoning tasks (Yang et al., 2025b). In the IT domain, MMMG introduces a novel dataset for multimodal knowledge graph generation, advancing knowledge-structured reasoning in technical contexts (Yao et al., 2025c). SciVer provides 3,000 expert-labeled examples for multimodal scientific claim verification across four distinct reasoning types derived from real research papers (Wang et al., 2025c).
@@ 5.4.1 GUI Navigation
Benchmarks in GUI navigation assess agents’ abilities to plan and execute tasks across digital interfaces, requiring robust visual-language grounding and multi-step reasoning. WebArena (Zhou et al., 2024c) and Mind2Web (Deng et al., 2023) offer realistic web environments for navigation and information extraction, with Mind2Web further introducing cross-website tasks to test generalizability. VisualWebBench (Liu et al., 2024b) advances visual-intensive planning with 1.5K tasks focused on cross-page integration and element localisation. Windows Agent Arena (Bonatti et al., 2024) evaluates cross-application planning in desktop environments, while Ferret-UI (You et al., 2024) focuses on grounded UI understanding for executing multi-step instructions. Benchmarks like WebShop (Yao et al., 2022) test visual-language grounding in simulated e-commerce environments. Similarly, OSWorld (Xie et al., 2024a) and OmniACT (Kapoor et al., 2024) provide real desktop OS environments, supporting cross-application workflows such as file manipulation and data processing. VisualAgentBench (Liu et al., 2024d) extends this paradigm by systematically evaluating large multimodal models across GUI, embodied, and visual design tasks, establishing a unified benchmark for planning and acting in visually rich digital environments. This is complemented by benchmarks like LlamaTouch (Zhang et al., 2024d), which scales mobile UI automation with 495 tasks, testing multi-step operations such as app navigation.
@@ 5.5 Evaluation Method
Due to the diversity of answers, exact and fuzzy matching methods are often unable to encompass all candidate options. To ensure fairness and accuracy in evaluation, the Options Matching approach was introduced. In this method, the system prompt includes several candidate options, and the model is required to select the most appropriate one. Moreover, to reduce the possibility of the model exhibiting a preference for a specific option during the selection process, works such as MMBench (Liu et al., 2024f) have adopted the CircularEval methodology to minimise stochastic variations in the evaluation.
@@
During the evaluation process, the capabilities of a single model are inherently limited, which may lead to shortcomings when processing diverse multimodal information. To this end, agent-based approaches can leverage tools to mitigate the inherent limitations of the model itself. For instance, CIGEval (Wang et al., 2025n) expands the visual understanding capabilities of MLLMs by integrating a multi-functional toolbox, thereby enabling more fine-grained evaluation. Moreover, multi-agent discussions have shown effectiveness in downstream tasks by fostering consensus and producing more robust solutions (Xu et al., 2023b; Chen et al., 2024d; Xu et al., 2025i), a benefit that also extends to evaluation settings. Methods that leverage cooperative or adversarial interactions among multiple agents to assess outputs have demonstrated more reliable and interpretable evaluations (Chan et al., 2024; Li et al., 2024h; Zhao et al., 2024b; Liang et al., 2024).
@@ 6 Conclusion
In this paper, we survey the evolution of multimodal reasoning models, highlighting pivotal advancements and paradigm-shifting milestones in the field. While current models predominantly adopt a␣language-centric reasoning paradigm, delivering impressive results in tasks like visual question answering, visual math, and video understanding. Notably,␣visual-centric long reasoning (e.g., understanding 3D contexts, addressing complex visual information-seeking questions) and␣interactive multimodal reasoning (e.g., dynamic cross-modal dialogue or iterative feedback loops) remain underdeveloped frontiers requiring deeper exploration.
Building on empirical evaluations and experimental insights, we propose a forward-looking concept for␣inherently large multimodal models that transcend language-dominated architectures. Such models should prioritize three core capabilities: Multimodal Agentic Reasoning: Enabling proactive environmental interaction (e.g., embodied AI agents that learn through real-world trial and error). Omini-Modal Understanding and Generative Reasoning: Integrating any-modal semantics (e.g., aligning abstract concepts across vision, audio, and text) while resolving ambiguities in complex, open-world contexts; Producing coherent, context-aware outputs across modalities (e.g., generating diagrams from spoken instructions or synthesizing video narratives from text). By addressing these dimensions, future models could achieve human-like contextual adaptability, bridging the gap between isolated task performance and generalized, real-world problem-solving.